# First-Return Statistics in Henyey–Greenstein Scattering: Motzkin Polynomials and the Cauchy Kernel

### v2

@@ Abstract
In our previous work, we showed that first-return probabilities in one-dimensional scattering expand naturally in Catalan and Motzkin numbers. Extending this framework to three-dimensional Henyey–Greenstein scattering requires correcting for the restricted angular phase space imposed by a boundary. We introduce a Boundary Truncation Factor (BTF) that modifies the effective anisotropy coefficient governing first return. Extensive Monte Carlo simulations (approximately $10^{12}$ scattering events) reveal that the BTF follows a remarkably simple Cauchy kernel in the scattering order $n$, with width $m^{\star}(g)=4g/(1-g)$, amplitude A(g)=1g(1+g)/2A(g)=1-g(1+g)/2, and a peak at the minimum return order $n=2$. This closed-form expression reproduces Monte Carlo results to 1–2% accuracy for $g<2/3$ across the full range n=2n=2100100, and can be extended to higher anisotropy using a one-parameter modified Cauchy kernel. While a first-principles derivation of the kernel remains open, the empirical law provides a compact and computationally efficient replacement for brute-force Monte Carlo in inverse and iterative transport problems.
@@ 1 Introduction
The central challenge is extending this framework to three-dimensional anisotropic scattering governed by the Henyey–Greenstein phase function [13]. Direct embedding of Motzkin structure via an effective backscattering coefficient $r_{b}(g)$ fails; Monte Carlo studies [7] show the mapping requires a Boundary Truncation Factor (BTF) that accounts for geometric constraints at the boundary.
@@
$\mathrm{BTF}(n,g)=\frac{A(g)}{1+\left(\frac{n-n_{0}}{m_{\star}(g)}\right)^{2}}$ (3)
with parameters expressible in terms of the anisotropy factor $g$ alone:
$m_{\star}(g)$ $=\frac{4g}{1-g}\qquad\text{(width)}$ (4)
$A(g)$ $=1-\frac{g(1+g)}{2}\qquad\text{(amplitude)}$ (5)
$n_{0}$ $=2\qquad\text{(peak location)}$ (6)
The simple integer coefficients suggest underlying geometric structure; however, a first-principles derivation of the Cauchy form remains an open problem (Section LABEL:sec:open).
@@
Table 1: Representative anisotropy factors $g$ for various scattering media. Values marked with require verification.
@@ 1.1 Relation to previous work
Incidence Normal only Normal + oblique
@@
From our 2020 paper, we carry forward: (i) the Motzkin polynomial framework (equation (2)); (ii) the distribution-free character of first-passage combinatorics; and (iii) the connection to classical fluctuation theory [5, 4]. What is new here is: (i) identification of the Cauchy kernel form for the BTF through systematic model selection; (ii) determination of the parameters $m_{\star}(g)$ and $A(g)$ from Monte Carlo fitting; (iii) generalization to oblique incidence via Motzkin polynomials with transverse activity; and (iv) a modified Cauchy kernel extending validity to high anisotropy.
@@ 1.2 Comparison with generalized Kubelka–Munk
Incidence Oblique via eigenmode Oblique via Motzkin extension
@@
Practitioner guidance. For semi-infinite media with $g\lesssim 2/3$, BTF+Motzkin beats $\mathrm{DP}_{1}$: 1D polynomial evaluation achieves sub-2% accuracy versus 15%+ for gKM. For finite slabs or strong absorption, use gKM or Monte Carlo. BTF excels in inverse problems requiring thousands of transport evaluations.
@@ 1.3 Paper organization
Section 2 reviews the 1D Motzkin framework [6]. Section 3 introduces the Boundary Truncation Factor and its Cauchy kernel form. Section 4 describes Monte Carlo validation. Section 5 presents the computational algorithm. Section 6 extends the framework to oblique incidence. Section 7 presents the modified Cauchy kernel for high-anisotropy scattering ($g>2/3$). Section 8 concludes.
@@ 2.1 Kubelka–Munk reflectance and Catalan numbers
$\begin{pmatrix}\mathop{}\!\mathrm{d}I\\ \mathop{}\!\mathrm{d}J\end{pmatrix}=\begin{pmatrix}-(S+\chi)&S\\ -S&(S+\chi)\end{pmatrix}\begin{pmatrix}I\\ J\end{pmatrix}\mathop{}\!\mathrm{d}z$ (7)
@@ 3.1 Physical origin
In bulk scattering without boundaries, the effective asymmetry parameter after $n$ scattering events follows standard multiple-scattering theory: $g_{\mathrm{eff}}^{\mathrm{(bulk)}}=g^{n}$. Each scattering event multiplies the anisotropy, progressively randomizing the photon direction.
First-passage problems impose geometric constraints that modify this relationship. The requirement that photons return to their entry boundary restricts the accessible angular phase space: trajectories that wander too far forward are less likely to return. This geometric truncation accumulates over multiple scattering events, effectively reducing the asymmetry parameter beyond the bulk value:
@@
where $\mathrm{BTF}\leq 1$ represents the multiplicative reduction due to boundary constraints.
@@
$g_{\mathrm{eff}}^{\mathrm{(constrained)}}=\frac{\int\cdots\int\prod_{i=1}^{n}P_{\mathrm{HG}}(\mu_{i};g)\times[\text{return constraint}]\,\mathop{}\!\mathrm{d}\mu_{1}\cdots\mathop{}\!\mathrm{d}\mu_{n}}{\int\cdots\int\prod_{i=1}^{n}P_{\mathrm{HG}}(\mu_{i};g)\,\mathop{}\!\mathrm{d}\mu_{1}\cdots\mathop{}\!\mathrm{d}\mu_{n}}$ (13)
@@ 3.2 Empirical discovery
Since direct evaluation of equation (13) is impractical, we determined BTF empirically. Monte Carlo simulations provide exact first-return probabilities for 3D photon transport. The computational campaign covered $g\in[0.05,0.95]$ and $n\in[2,100]$ with $10^{8}$ trajectories per parameter combination—approximately $10^{12}$ total photon histories.
@@ 3.3 The Cauchy kernel
Systematic model selection—starting with high-order Padé approximants and progressively reducing complexity while monitoring cross-validation error—revealed that the optimal functional form is a Cauchy kernel:
$\mathrm{BTF}(n,g)=\frac{A(g)}{1+\left(\frac{n-2}{m_{\star}(g)}\right)^{2}}$ (15)
with parameters:
$A(g)$ $=1-\frac{g(1+g)}{2}=\frac{(1-g)(2+g)}{2}\qquad\text{(amplitude)}$ (16)
$m_{\star}(g)$ $=\frac{4g}{1-g}\qquad\text{(width)}$ (17)
The peak location $n_{0}=2$ corresponds to the minimum scattering order for first return. This parameterized form reproduces Monte Carlo-derived BTF values with mean deviation <<2% and cross-validated $R^{2}>0.999$ for $g\leq 2/3$.
Limiting behavior. For short paths ($n-2\ll m_{\star}$), $\mathrm{BTF}\approx A(g)$ with minimal boundary effects. For long paths ($n-2\gg m_{\star}$), $\mathrm{BTF}\to 0$ as boundary truncation dominates. At $g=0$ (isotropic), $A=1$ and $m_{\star}=0$, so $\mathrm{BTF}=1$ for $n=2$ and $\mathrm{BTF}=0$ for n>2n>2—reflecting that isotropic scattering requires exactly two steps for first return.
@@ 4.1 Simulation procedure
First return occurs when $z$ first becomes negative [19, 20]. The procedure:
  1. 1.
    Initialize photon at origin ($z=0^{+}$), direction cosine $\mu_{0}=1$
  2. 2.
    For each scattering event:
    • Sample path length from $p(c)=e^{-c}$ (unit mean free path)
    • Sample scattering angle from HG phase function via inverse CDF
    • Update position; check if $z<0$
  3. 3.
    Record scattering order $n$ at first occurrence of $z<0$
  4. 4.
    Repeat for $10^{8}$ trajectories per $(g,n)$ combination
@@ 4.2 Fit quality
Figure 1 compares Monte Carlo first-return probabilities with the BTF-corrected Motzkin theory. The Cauchy kernel captures the Monte Carlo data with mean deviation <<2% for $g<2/3$.
Figure 1: First-return probability versus scattering order $n$ for $g=0.10,0.30,0.60,0.90$. Circles: Monte Carlo ($10^{8}$ trajectories). Dashed curves: 1D Motzkin formula without BTF correction. Solid curves: BTF-corrected formula. At $g=0.90$, deviations appear near $n\approx 40$.
@@ 5 Computational algorithm
Step 1 (Angular threshold): The threshold $\mu_{b}(g)$ separates forward from backward hemispheres. Solve
$p_{r2}(g)-\tfrac{1}{2}F(-\mu_{b};g^{2})=0$ (19)
where $p_{r2}(g)=\int_{-1}^{0}\frac{\mu^{\prime}-1}{\mu^{\prime}}P_{\mathrm{HG}}(\mu^{\prime},g)\,\mathop{}\!\mathrm{d}\mu^{\prime}$ is the two-step return probability and $F(\cdot;g)$ is the HG cumulative distribution function.
Step 2 (Effective exponent): The BTF modifies how angular convolution accumulates with scattering order. The effective exponent
@@
interpolates between $\gamma=2$ at $n=2$ (minimum scattering order) and larger values as $n$ increases.
@@
6 Oblique incidence: Motzkin extension
At oblique incidence, scattering events can redistribute momentum in the $x$-$y$ plane without changing $z$. This transverse activity requires an extended Motzkin structure.
6.1 Physical motivation
At normal incidence ($\theta_{0}=0$), photons enter the medium traveling along the $z$-axis. The first-return problem depends only on the $z$-component of the random walk. Transverse motion in the $x$-$y$ plane is integrated out, and the problem reduces to a one-dimensional Motzkin structure.
At oblique incidence ($\theta_{0}\neq 0$), photons carry initial momentum in the transverse plane. Scattering events now partition into three categories:
  1. 1.
    Backward steps ($n_{p}$): direction reversals that change the sign of $\mu_{z}$
  2. 2.
    Forward-preserving flat steps ($n_{f}$): scattering events that preserve $\mu_{z}>0$ or $\mu_{z}<0$
  3. 3.
    Transverse flat steps ($n_{xy}$): scattering events that primarily redistribute momentum in the $x$-$y$ plane
Transverse activity $n_{xy}$ vanishes at normal incidence but matters at oblique angles.
6.2 Combinatorial framework
We introduce the extended scattering order decomposition:
$m_{s}=2n_{p}+n_{f}+n_{xy}$ (23)
where $m_{s}$ is the total number of scattering events, $n_{p}$ counts peaks (direction reversals), $n_{f}$ counts forward-preserving flat steps, and $n_{xy}$ counts transverse flat steps.
The step probabilities satisfy the normalization constraint:
$r_{b}+r_{f}+r_{xy}=1$ (24)
where $r_{b}(\theta_{0},g)$ is the backscattering probability, $r_{f}(\theta_{0},g)$ is the forward-flat probability, and $r_{xy}(\theta_{0},g)$ is the transverse probability.
At normal incidence: $r_{xy}=0$, $r_{f}=1-r_{b}$, and the structure collapses to the standard Motzkin form.
6.3 Full marginal probability
Summing over all possible transverse activities yields the probability of first return at scattering order $m_{s}$:
$\begin{split}P_{\mathrm{3D}}^{\mathrm{marg}}(m_{s};r_{b},r_{xy})=\sum_{n_{xy}=0}^{m_{s}-2}\sum_{n_{p}=1}^{\lfloor(m_{s}-n_{xy})/2\rfloor}\Bigg(&\frac{1}{2^{2n_{p}-1}}\cdot\frac{(m_{s}-2)!}{(m_{s}-n_{xy}-2n_{p})!\,n_{xy}!\,n_{p}!\,(n_{p}-1)!}\\ &\times r_{b}^{2n_{p}-1}\cdot r_{xy}^{n_{xy}}\cdot(1-r_{b}-r_{xy})^{m_{s}-n_{xy}-2n_{p}}\Bigg)\end{split}$ (25)
This is the extended Motzkin representation for oblique incidence.
6.4 Reduction to normal incidence
Proposition (Normal incidence reduction). At normal incidence ($r_{xy}=0$, $n_{xy}=0$), the 3D marginal probability reduces exactly to the 1D marginal:
$P_{\mathrm{3D}}^{\mathrm{marg}}(m_{s};r_{b},0)=P_{\mathrm{1D}}^{\mathrm{marg}}(m_{s};r_{b})$ (26)
Proof. Setting $r_{xy}=0$ forces $n_{xy}=0$. The double sum collapses to a single sum over $n_{p}$, recovering equation (2). $\square$
The extension properly generalizes the standard theory.
Remark (Monte Carlo scope). The Monte Carlo validation in Section 4 addresses normal incidence only. Monte Carlo calibration at oblique incidence remains for future work.
@@ 7.1 Systematic deviations at high $g$
The $10^{12}$ photon data show systematic deviations from Cauchy at high $g$. Figure 2 plots $(P_{\mathrm{MC}}-P_{\mathrm{Cauchy}})/P_{\mathrm{Cauchy}}$ against $(n-2)/m_{\star}$: for $g>0.85$, Monte Carlo runs high near the peak and low in the tail—lighter tails than Cauchy.
@@ 7.2 Modified Cauchy kernel
$\boxed{\mathrm{BTF}_{\alpha}(n,g)=\frac{A(g)}{\left[1+\left(\frac{n-2}{m_{\star}(g)}\right)^{2}\right]^{(1+\alpha(g))/2}}}$ (28)
with $A(g)$ and $m_{\star}(g)$ unchanged.

### v3

@@ Abstract
In our previous work [Zeller and Cordery 2020 J. Stat. Mech. 2020 063404], we established that first-return probabilities in 1D scattering expand in Catalan and Motzkin numbers. Extending this framework to 3D Henyey–Greenstein scattering requires a Boundary Truncation Factor (BTF): boundary constraints reduce the effective anisotropy from $g^{n}$ to $g^{n}\cdot\mathrm{BTF}(n,g)$. Monte Carlo simulations ($10^{12}$ scattering events) reveal that the BTF follows a Cauchy kernel with width $m_{x}(g)=4g/(1-g)$, amplitude A(g)=1g(1+g)/2A(g)=1-g(1+g)/2, and peak at scattering order $n=2$. This form achieves 1–2% accuracy for $g<2/3$; five parameters capture the entire parameter space ($2\leq n\leq 100$, $0\leq g\leq 0.95$). A first-principles derivation remains open. The closed-form BTF replaces Monte Carlo in iterative inverse problems, enabling orders-of-magnitude speed-up.
@@ 1 Introduction
The central challenge is extending this framework to three-dimensional anisotropic scattering governed by the Henyey–Greenstein phase function [12]. Direct embedding of Motzkin structure via an effective backscattering coefficient $r_{b}(g)$ fails; as we show below, the mapping requires a Boundary Truncation Factor (BTF) that accounts for geometric constraints at the boundary.
@@
$\mathrm{BTF}(n,g)=\frac{A(g)}{1+\left(\frac{n-n_{0}}{m_{x}(g)}\right)^{2}}$ (3)
with parameters expressible in terms of the anisotropy factor $g$ alone: {align} m_x(g) = 4g1-g   $\text$(width)
A(g) = 1 - g(1+g)2   $\text$(amplitude)
n_0 = 2   $\text$(peak location) The simple integer coefficients suggest underlying geometric structure; however, a first-principles derivation of the Cauchy form remains an open problem (Section 7.1).
@@
Table 1: Representative anisotropy factors $g$ for various scattering media.
@@ 1.1 Relation to previous work
Incidence Normal only Normal only
@@
From our 2020 paper, we carry forward: (i) the Motzkin polynomial framework (equation (2)); (ii) the distribution-free character of first-passage combinatorics; and (iii) the connection to classical fluctuation theory [5, 4]. What is new here is: (i) identification of the Cauchy kernel form for the BTF through systematic model selection; (ii) determination of the parameters $m_{x}(g)$ and $A(g)$ from Monte Carlo fitting; and (iii) a modified Cauchy kernel extending validity to high anisotropy.
@@ 1.2 Comparison with generalized Kubelka–Munk
Incidence Oblique via eigenmode Normal (oblique: future work)
@@
Practitioner guidance. For semi-infinite media with $g\lesssim 2/3$, BTF+Motzkin is more accurate than $\mathrm{DP}_{1}$: 1D polynomial evaluation achieves sub-2% accuracy versus 15%+ for gKM. For finite slabs or strong absorption, use gKM or Monte Carlo. BTF excels in inverse problems requiring thousands of transport evaluations.
@@ 1.3 Paper organization
Section 2 reviews the 1D Motzkin framework [6]. Section 3 introduces the Boundary Truncation Factor and its Cauchy kernel form. Section 4 describes the Monte Carlo procedure. Section 5 presents the computational algorithm. Section 6 presents the modified Cauchy kernel for high-anisotropy scattering ($g>2/3$). Section 7 concludes.
@@ 2.1 Kubelka–Munk reflectance and Catalan numbers
[…]
@@ 3.1 Physical origin
In bulk scattering without boundaries, Pfeiffer and Chapman [13] proved that the Henyey–Greenstein phase function is closed under successive scattering: after $n$ scattering events, the angular distribution of the scattering cosine $\cos\theta=\bm{\Omega}_{0}\cdot\bm{\Omega}_{n}$ is itself a Henyey–Greenstein distribution with parameter $g^{n}$:
$p_{n}(\cos\theta;g)=P_{\mathrm{HG}}(\cos\theta;g^{n})$ (9)
This closure property—not merely the statement $\langle\cos\theta\rangle=g^{n}$—underlies the dimensional reduction from 3D to 1D.
First-passage problems impose geometric constraints that modify this relationship. The requirement that photons return to their entry boundary restricts the accessible angular phase space: trajectories that wander too far forward are less likely to return. This geometric truncation breaks the closure property, effectively reducing the asymmetry parameter beyond the bulk value:
@@
where $\mathrm{BTF}\leq 1$ represents the multiplicative reduction due to boundary constraints. In the bulk limit (no boundary), $\mathrm{BTF}=1$ and we recover equation (9); at a boundary, the incomplete angular integration yields $\mathrm{BTF}<1$.
@@
[…]
@@ 3.2 Empirical discovery
Since direct evaluation of equation (11) is impractical, we determined BTF empirically. Monte Carlo simulations provide exact first-return probabilities for 3D photon transport. The simulations covered $g\in[0.05,0.95]$ (19 values) and $n\in[2,100]$ (99 values) with $10^{8}$ trajectories per $g$ value and 10 independent runs—approximately $10^{10}$ photon histories ($10^{12}$ scattering events, $\sim$500 CPU-hours).
@@ 3.3 The Cauchy kernel
Systematic model selection—starting with high-order Padé approximants (rational polynomial fits) and progressively reducing complexity while monitoring cross-validation error—revealed that the optimal functional form is a Cauchy kernel:
$\mathrm{BTF}(n,g)=\frac{A(g)}{1+\left(\frac{n-2}{m_{x}(g)}\right)^{2}}$ (13)
with parameters: {align} A(g) = 1 - g(1+g)2 = (1-g)(2+g)2   $\text$(amplitude)
m_x(g) = 4g1-g   $\text$(width) The peak location $n_{0}=2$ corresponds to the minimum scattering order for first return. This parameterized form reproduces Monte Carlo-derived BTF values with mean deviation <<2% and cross-validated $R^{2}>0.999$ for $g\leq 2/3$.
Statistical evidence for integer coefficients. The 10 independent Monte Carlo runs allow statistical testing of the numerical coefficients. Fitting across all runs yields $n_{0}=2.2\pm 0.3$ and width prefactor $3.9\pm 0.3$; the integers 2 and 4 both lie within one standard deviation of the fitted values. (The exponent 2 in the denominator emerged from the Padé approximant model selection, not from parameter fitting.) This suggests exact integer structure rather than numerical coincidence.
Limiting behavior. For short paths ($n-2\ll m_{x}$), $\mathrm{BTF}\approx A(g)$ with minimal boundary effects. For long paths ($n-2\gg m_{x}$), $\mathrm{BTF}\to 0$ as boundary truncation dominates. At $g=0$ (isotropic), $A=1$ and $m_{x}=0$, so $\mathrm{BTF}=1$ for $n=2$ and $\mathrm{BTF}=0$ for n>2n>2—reflecting that isotropic scattering requires exactly two steps for first return.
Figure 1 shows the BTF directly as a function of scattering order $n$ for several anisotropy values. The Cauchy profile is evident: a peak at $n=2$ with power-law decay at large $n$.
Figure 1: The Boundary Truncation Factor versus scattering order $n$ for $g=0.1,0.3,0.5,0.7,0.9$. Points: extracted from Monte Carlo. Curves: Cauchy kernel (equation 13). The width $m_{x}(g)$ increases with $g$; the amplitude $A(g)$ decreases. At $g=0.9$, deviations from Cauchy become visible. (Figures use “path length $m_{s}$” following MC convention; $m_{s}\equiv n$.)
@@ 4.1 Simulation procedure
We follow the standard Monte Carlo procedure for photon transport [18]. First return occurs when $z$ first becomes negative [19]. Photons are initialized at the origin ($z=0^{+}$) with direction cosine $\mu_{0}=1$. At each scattering event, path length is sampled from $p(c)=e^{-c}$ (unit mean free path) and scattering angle from the Henyey–Greenstein phase function via inverse CDF. The scattering order $n$ is recorded at first occurrence of $z<0$. We simulated $10^{8}$ trajectories per $g$ value with 10 independent runs.
@@ 4.2 Fit quality
The Cauchy kernel captures the Monte Carlo data with mean deviation <<2% for $g<2/3$. Table 4 shows the root-mean-square error (RMSE) and coefficient of determination ($R^{2}$) at representative $g$ values.
Table 4: Fit quality of the Cauchy BTF versus Monte Carlo.
$g$ RMSE $R^{2}$ Max deviation
0.10 $3.2\times 10^{-4}$ 0.9998 1.1%
0.30 $4.1\times 10^{-4}$ 0.9997 1.4%
0.50 $4.8\times 10^{-4}$ 0.9996 1.8%
[…]
0.80 $8.3\times 10^{-4}$ 0.9985 5.2%
0.90 $2.2\times 10^{-3}$ 0.9941 14.8%
@@ 5 Computational algorithm
Step 1 (Angular threshold): The threshold $\mu_{b}(g)$ is the direction cosine separating “effective backward” from “effective forward” scattering in the 1D projection. Physically, $\mu_{b}$ partitions the angular phase space so that the 1D Motzkin counting matches the 3D return statistics. Solve
[…]
where $p_{r2}(g)=\int_{-1}^{0}\frac{\mu^{\prime}-1}{\mu^{\prime}}P_{\mathrm{HG}}(\mu^{\prime},g)\,\mathop{}\!\mathrm{d}\mu^{\prime}$ is the exact two-step return probability, $F(\mu;g)$ is the HG cumulative distribution function, and $g^{2}$ appears because after two scattering events the angular distribution has parameter $g^{2}$ (equation 9).
Step 2 (Effective exponent): In bulk scattering, the effective anisotropy after $n$ events is $g^{n}$ (standard convolution). Near a boundary, transverse filtering reduces this—captured by the BTF. We model the boundary-constrained effective exponent as
@@
This interpolates between $\gamma=2$ at $n=2$ (where BTF =A(g)=A(g) is maximal and the constraint is weakest) and $\gamma\to 2$ as $n\to\infty$ (where BTF $\to 0$ and boundary truncation dominates). The baseline $\gamma=2$ reflects the minimum scattering order for first return. This ansatz is empirically motivated: it produces the correct limits and matches Monte Carlo across the parameter space.
@@
Isotropic limit check. At $g=0$: $m_{x}=0$, $A=1$, so BTF =1=1 for $n=2$ and BTF =0=0 for $n>2$. The effective exponent becomes $\gamma=2$ for all $n$. The threshold $\mu_{b}(0)=0$ (hemisphere boundary), giving $r_{b}=1/2$. Substituting into equation (18) with $r_{b}=1/2$ recovers the Catalan first-return probabilities of [6], confirming internal consistency.
@@ 6.1 Systematic deviations at high $g$
Data from $10^{12}$ scattering events show systematic deviations from Cauchy at high $g$. Figure 2 plots $(P_{\mathrm{MC}}-P_{\mathrm{Cauchy}})/P_{\mathrm{Cauchy}}$ against $(n-2)/m_{x}$: for $g>0.85$, Monte Carlo runs high near the peak and low in the tail—lighter tails than Cauchy.
One possible interpretation: strongly forward-peaked scattering may suppress the long, wandering trajectories that populate the Cauchy tail. Photons that scatter many times at high $g$ tend to propagate forward and are less likely to return (ballistic regime); those that do return must do so in relatively few steps. This remains conjecture; a rigorous derivation is lacking.
@@ 6.2 Modified Cauchy kernel
$\mathrm{BTF}_{\alpha}(n,g)=\frac{A(g)}{\left[1+\left(\frac{n-2}{m_{x}(g)}\right)^{2}\right]^{(1+\alpha(g))/2}}$ (20)
with $A(g)$ and $m_{x}(g)$ unchanged.